Round the following list of numbers to three significant figures.

a. 10,000

b. .00321

c. 0.397000

d. 15.9

e. 0.0421


1
Expert's answer
2020-12-01T09:12:39-0500

Number three significant figures


a. 10,000 = 10.0×10^3


b. .00321 = 3.21×10^-3


c. 0.397000 = 3.97×10^-1


d. 15.9 = 15.9


e. 0.0421 = 4.21×10^-2